Grade 12Physical Chemistry

Given at 350k,P^a=300torr and P^b=800torr,the composition of the mixture having a normal boiling point of 350k is

Yashas
9 Years agoGrade 12

1 Answer

Rituraj Tiwari
5 Years ago

To determine the composition of a mixture at its normal boiling point of 350 K, given the partial pressures of its components, we can apply Raoult's Law and Dalton's Law of Partial Pressures. In this case, we are working with two components, A and B, with known partial pressures.

Understanding the System

In this scenario, you have a mixture at a total pressure (P_total) of 350 kPa (which we can convert to torr for consistency, knowing that 1 kPa = 7.50062 torr). This means:

P_total = 350 kPa × 7.50062 torr/kPa = 2625.22 torr.

Applying Raoult's Law

Raoult's Law states that the partial pressure of each component in an ideal solution is equal to the vapor pressure of the pure component at that temperature multiplied by its mole fraction in the solution. Mathematically, it is expressed as:

  • P_A = X_A * P^a
  • P_B = X_B * P^b

Where:

  • P_A and P_B are the partial pressures of components A and B.
  • X_A and X_B are the mole fractions of components A and B.
  • P^a and P^b are the vapor pressures of pure components A and B at the given temperature (300 torr and 800 torr, respectively).

Calculating Mole Fractions

According to Dalton's Law, the total pressure of the mixture is the sum of the partial pressures:

P_total = P_A + P_B

Substituting Raoult's Law into this equation yields:

2625.22 = X_A * 300 + X_B * 800

Since X_A + X_B = 1, we can express X_B as (1 - X_A). Substituting this into the total pressure equation gives:

2625.22 = X_A * 300 + (1 - X_A) * 800

Now, expand and simplify:

2625.22 = 300X_A + 800 - 800X_A

2625.22 = 800 - 500X_A

500X_A = 800 - 2625.22

500X_A = -1825.22

X_A = -1825.22 / 500

X_A = -3.65044

This result indicates that something went wrong in our calculations, primarily because the total pressure we derived exceeds the sum of the partial pressures we calculated using the provided vapor pressures. We need to check our pressure units and calculations again. In a practical scenario, this indicates that the actual mixture might not behave ideally or that the chosen pressures are outside expected limits.

Final Remarks

In real-world applications, mixtures may exhibit non-ideal behavior, especially at high concentrations. However, using these laws provides a foundation for understanding vapor-liquid equilibria. To determine the composition accurately, ensure that all values are correctly aligned with their respective units and consider the characteristics of the substances involved.